OpenStudy (anonymous):

Which of the following is equivalent to ab = 7? Answer logab = 7 logb7 = a logba = 7 loga7 = b

OpenStudy (anonymous):

loga7 = b coorrrrreeect?

OpenStudy (anonymous):

Well, for any of these to work, we have a notational problem. What you meant to say was\[a^b = 7\]Keeping in mind the most important idea in the whole chapter, which is\[a^x=y \iff \log_{a} y=x\]We can substitute to get\[a^b=7 \iff \log_a7=b\]

OpenStudy (anonymous):

If we overlook the notation problems, you have the right idea. Good job.

OpenStudy (anonymous):

thank youuuu :]

OpenStudy (anonymous):

No sweat.